# Machine Space I: Weak exponentials and quantification over compact spaces

### v3

@@ 3. Machine space
A point $x\in X$ is accepted by a machine $m$ if there exists some $J_{k}$ such that $x$ is accepted by each ‘basic machine’ $g_{j}$ for all $j\in J_{k}$ (i.e. such that $x$ lies in the open given by $g_{j}$ for all such $j$). In this way, each machine $m$ will define an open on which it halts. Many of these machines will correspond to the same opens, as described by the relation $R$, so that each open in $X$ corresponds to an equivalence class of machines (see Definition 3.7 and Proposition 3.8).
In what follows we will formally define the space of machines and relate it to the exponential $\Sigma^{X}$, when it exists, and to the Scott topology on $\mathcal{O}X$ more generally.
3.1. Presentations and weak exponentials
For a general space $X$ the space of opens $\Sigma^{X}$ need not exist. However, given a presentation $\mathcal{O}X=\langle G\mid R\rangle$ there is however always a natural space of machines, whose points are precisely the machines described above and whose opens have a natural interpretation in terms of observable properties of machines. We claim that this space is $\Sigma^{\Sigma^{G}}$ and will justify this assertion in the paragraphs below.
Recall that $\langle G\mid R\rangle$ is a quotient of the free frame on $G$, which in turn is the frame of opens of $\Sigma^{G}$. The opens of $\Sigma^{G}$ are precisely the machines mentioned in Definition 3.1.
The frame quotient $\Sigma^{G}\twoheadrightarrow\langle G\mid R\rangle$ exhibits $X$ as a sublocale of $\Sigma^{G}$. The points of $\Sigma^{G}$ correspond to subsets of $G$ and we can then interpret this inclusion as sending points of $X$ to the set of generating opens in which they lie. The points in the image of this inclusion correspond to the frame homomorphisms $\mathcal{O}(\Sigma^{G})\to\mathcal{O}1$ which respect the relations. General points of $\Sigma^{G}$ may then be interpreted as ‘generalised points’ of $X$ and can be given as input to machines on $X$. This perspective will be useful when we study compactness in Section 4.
So machines on $X$ correspond to opens of $\Sigma^{G}$. But $\Sigma^{G}$ is always locally compact and so we can consider $\Sigma^{\Sigma^{G}}$ — the space of opens of $\Sigma^{G}$. This will be our space of machines for $X$. The topology on $\Sigma^{\Sigma^{G}}$ is that of the double powerlocale on $G$ (see [29, 31]) and is generated by opens ${\boxtimes}U$ for each $U\subseteq G$ where a machine $m=\bigvee_{i\in I}\bigwedge_{j\in J_{i}}g_{j}$ lies in ${\boxtimes}U$ if and only if there exists some $J_{k}$ such that $g_{j}\in U$ for each $j\in J_{k}$. Intuitively this means we can semi-decide if there exists some parallel branch of a machine $m$ such that each of the generators that run in that branch are contained in $U$.
The fact that we have nontrivial verifiable properties indicates that these machines are not complete black boxes and that there is some information that leaks while they run. In order to understand this better, let us consider the following compelling but flawed model of the situation, before showing how to remedy it. Note that with all such “real-world” models the set of generators is assumed to be countable.
Model 3.2.
Let $X$ be a locale with $\mathcal{O}X=\langle G\mid R\rangle$. We take each generator $g\in G$ to be a black box which accepts points of $X$ as input and either halts after some finite time or runs forever. We imagine these generators arranged in a line on some suitably long table. Also on the table are a number of mobile robots which themselves accept points of $X$ as input and will roll from generator to generator testing the point in the following manner.
The robot progresses through different stages. On the $i^{\text{th}}$ stage it selects a group of finitely many generators, writes the number $i$ onto them and (if they are not already running) runs them with the point as input. The generators are left to run as the next stage is started. If at any point every generator in a group has halted, the a light on the robot’s head flashes green indicating that it accepts the point.
Such a robot instantiates a machine $\bigvee_{i\in I}\bigwedge_{j\in J_{i}}g_{j}$ where each $J_{i}$ is the finite set of generators selected at stage $i$. ∎
If we take these robots to be our machines, then we can indeed verify ${\boxtimes}U$ by simply checking that every machine in some group is contained in $U$. However, there are additional questions that we can semi-decide which do not occur as opens in $\Sigma^{\Sigma^{G}}$, such as:
  1. (1)
    “is the generator $g_{1}$ visited before the generator $g_{2}$?”
  2. (2)
    “is the generator $g$ ever visited?”
  3. (3)
    “is the group $\{g_{1},g_{2}\}$ selected at least twice?”
The problem is that the robots, as described, do not actually correspond to points of $\Sigma^{\Sigma^{G}}$, rather they are points in a ‘freer’ structure in which $\wedge$ and $\vee$ are not assumed to be commutative, idempotent or absorbent. We can eliminate questions of the first kind by making the robot choose the order of the generators visited (both within and between groups) nondeterministically. For questions of the second kind, note that the machine $g$ should be identified with the machine $g\vee(g\wedge h)$. So if the robot visits the generators $S\subseteq G$ in a particular branch, then we should require it to visit a branch with generators $T$ for all $T\supseteq S$. This will ensure that questions of the second form will always be true and hence trivially verifiable. Finally, we ensure that no group of machines is selected more than once to render third question trivial in a similar way.
Note that original model already captured associativity and distributivity of the operations. Distributivity follows the fact that the formal operations are already reduced the canonical form of a join of meets. These axioms together are then enough to give a frame, which ensures that no extraneous verifiable properties of this sort exist.
Of course, if actually implemented in the real world there would almost certainly be additional questions one could ask pertaining to the specific engineering of the robots, manufacturing defects, and so on. However, we consider an idealised setting where the only information that can be gleaned is related to the robot’s interaction with the generators.
We believe this idealised model is still compelling evidence that such a space of machines is realisable. Thus, the space of machines is a sense more concrete than the space of opens, even when the latter exists, and it is interesting to see how we might obtain the space of opens from machine space. We will do this in Section 3.2.
Notice that $\Sigma^{\Sigma^{G}}$ by itself contains no information about the behaviour of the machines on inputs from the locale $X$ with presentation $\langle G\mid R\rangle$, as this data is contained in the relation $R$. We can represent this data if we equip $\Sigma^{\Sigma^{G}}$ with an evaluation map $\widetilde{\mathrm{ev}}\colon\Sigma^{\Sigma^{G}}\times X\to\Sigma$ which records whether a point lies in a machine. This makes $\Sigma^{\Sigma^{G}}$ a weak exponential as we show below. (Compare the construction of weak exponentials in $\mathrm{Top}$ given in [21].)
Definition 3.3.
@@ 3.1. Presentations and weak exponentials
Let $X$ and $B$ be objects in some category. We say $Z$ is a weak exponential for $B^{X}$ if there exists a morphism $\widetilde{\mathrm{ev}}\colon Z\times X\to B$ such for any $f\colon Y\times X\to B$ there exists a not necessarily unique $\overline{f}\colon Y\to Z$ such that the following diagram commutes.
{tikzpicture}
@@
{tikzpicture}
[…]
Observe that $u$ defines an open in $A\times X$. Since $A\times X$ is a subspace of $A\times\Sigma^{G}$, there exists a (not necessarily unique) open $u^{\prime}$ in $A\times\Sigma^{G}$ which restricts to give $u$. By the universal property of the exponential $\Sigma^{\Sigma^{G}}$, we obtain a morphism $\overline{u}^{\prime}$ making the bottom right triangle commute. It is not hard to see that the left-hand trapezium commutes and then a simple diagram chase confirms that $u=\widetilde{\mathrm{ev}}\circ(\overline{u}^{\prime}\times X)$. ∎
We arrive at the following definition.
@@
Let $X$ be a locale with $\mathcal{O}X=\langle G\mid R\rangle$. We define the associated (idealised) machine space to be $\Sigma^{\Sigma^{G}}$ together with the evaluation map $\widetilde{\mathrm{ev}}\colon\Sigma^{\Sigma^{G}}\times X\to\Sigma$ defined in Proposition 3.4. ∎
@@
In fact, for many purposes any weak exponential for $\Sigma^{X}$ behaves like a machine space for $X$. This could be interesting to explore further. However, for this paper we restrict our consideration to $\Sigma^{\Sigma^{G}}$, since the compactness algorithm given in Section 4 is dependent on the presentation of the locale by generators and relations.
@@ 3.2. Relationship between machine space and the space of opens
{tikzpicture}
[…]
@@
Let $X$ be a locally compact locale with $\mathcal{O}X=\langle G\mid R\rangle$. The map $q\colon\Sigma^{\Sigma^{G}}\to\Sigma^{X}$ as defined in Definition 3.7 has a section $s$ satisfying $\widetilde{\mathrm{ev}}\circ(s\times X)=\mathrm{ev}$.
@@
{tikzpicture}
[…]
It is immediate that $\widetilde{\mathrm{ev}}\circ(s\times X)=\mathrm{ev}$ and so all that remains is to prove that is is a section of $q$.
By the diagram we have $\mathrm{ev}\circ(q\times X)\circ(s\times X)=\widetilde{\mathrm{ev}}\circ(s\times X)=\mathrm{ev}=\mathrm{ev}\circ(\mathrm{id}_{\Sigma^{X}}\times X)$. Hence by the uniqueness condition of the universal property of the ‘true’ exponential $\Sigma^{X}$, we have $qs=\mathrm{id}_{\Sigma^{X}}$, as required. ∎
@@
The relationship between machine space and the exponential $\Sigma^{X}$ may be fruitfully interpreted from the perspective of verifiability. If one is not careful, one can be led to believe that the verifiable interpretation of topology implies that the space of opens $\Sigma^{X}$ should exist for all spaces $X$. This deficient reasoning would go as follows: the collection of opens should have some topological structure corresponding to the verifiable properties on it. Let us call this space $\mathcal{O}X$. Given an open $u\in\mathcal{O}X$ and a point $x\in X$, since $u$ is a verifiable property, we can verify whether $x$ lies in $u$. Carrying out this procedure should take a value in $\Sigma$ corresponding to whether the process ran forever or halted. Thus, this procedure represents a physically realisable evaluation map from $\mathcal{O}X\times X$ to $\Sigma$ which implies that this evaluation map is continuous. It is then not hard to check this would satisfy the properties of an exponential object, proving that $\mathcal{O}X=\Sigma^{X}$.
From the perspective of machine space the error lies in assuming that given an open and a point, it is possible to semi-decide membership of the point in the open. An open merely describes the verifiable property under consideration — it is a machine that actually physically carries out this verification. From this point of view, the fact that a section of the quotient map exists whenever $X$ is locally compact perfectly explains why a continuous evaluation map exists for $\mathcal{O}X$. Given an open $u$ and a point $x$, one can apply the section to $u$ to acquire a machine $s(u)$ which can then verify membership of $x$ in $u$.
On the other hand, when $X$ is not locally compact the argument breaks down and we will see that no way to continuously assign opens to machines in this case. Formulating the quotient of machine space for non-locally-compact locales presents some difficulties in the pointfree setting, so in the following section we will examine this from the point of view of point-set topology. In Section 3.2.2 we formalise what we can in the pointfree setting. The first section is sufficient for understanding the intuition, while the second is more technical.
3.2.1. The space of opens as a topological space
Recall that if a locale $X$ is given by $\mathcal{O}X=\langle G\mid R\rangle$ then the frame $\mathcal{O}X$ is a quotient of $\mathcal{O}(\Sigma^{G})$ by the congruence generated from the relations $R$. Let us call this frame quotient map $\overline{q}$. Note that $\mathcal{O}(\Sigma^{G})$ equipped with the Scott topology is precisely the machine space $\Sigma^{\Sigma^{G}}$. We can now use $\overline{q}$ to induce a topology on $\mathcal{O}X$.
Proposition 3.9.
@@ 3.2.1. The space of opens as a topological space
Proof.
We need to show that $U\subseteq\mathcal{O}X$ is Scott-open if and only if $\overline{q}^{-1}(U)$ is Scott-open. First suppose $U$ is Scott-open. We start by showing that $\overline{q}^{-1}(U)$ is an upset. Take $a\in\overline{q}^{-1}(U)$ and suppose $a\leq b$. Then $\overline{q}(a)\leq\overline{q}(b)$ and $\overline{q}(a)\in U$, giving $\overline{q}(b)\in U$. Hence, $b\in\overline{q}^{-1}(U)$ and $\overline{q}^{-1}(U)$ is an upset. Now suppose $\operatorname*{\bigvee{}^{\!\uparrow}}D\in\overline{q}^{-1}(U)$. This means $\overline{q}(D)\subseteq U$ and $\operatorname*{\bigvee{}^{\!\uparrow}}\overline{q}(D)\in U$. But $\overline{q}(D)$ is also directed and so, by Scott-openness of $U$, there is a $d\in D$ such that $\overline{q}(d)\in U$. Hence, $d\in\overline{q}^{-1}(U)$ and $\overline{q}^{-1}(U)$ is Scott-open.
Now take $U\subseteq\mathcal{O}X$ and suppose $\overline{q}^{-1}(U)$ is Scott-open. Say $a^{\prime}\in U$ and $a^{\prime}\leq b^{\prime}$. Now take $a$ to be any element of $\overline{q}^{-1}(a^{\prime})$ and $b$ to be an element of $\overline{q}^{-1}(b^{\prime})$ lying above $a$. This is always possible by joining with $a$ as necessary. Then $a\in\overline{q}^{-1}(U)$, as $a^{\prime}=\overline{q}(a)\in U$. Hence, $b\in\overline{q}^{-1}(U)$, since $\overline{q}^{-1}(U)$ is upward closed. Thus, $b^{\prime}=\overline{q}(b)\in U$ and so $U$ is an upset. Finally, suppose $\operatorname*{\bigvee{}^{\!\uparrow}}D^{\prime}\in U$. Set $D=\overline{q}^{-1}(D^{\prime})$. Note that $D$ is directed, that $\overline{q}(D)=D^{\prime}$ and that $\overline{q}(\operatorname*{\bigvee{}^{\!\uparrow}}D)=\operatorname*{\bigvee{}^{\!\uparrow}}\overline{q}(D)=\operatorname*{\bigvee{}^{\!\uparrow}}D^{\prime}\in U$. Thus, $\operatorname*{\bigvee{}^{\!\uparrow}}D\in\overline{q}^{-1}(U)$ and so there is a $d\in D$ such that $d\in\overline{q}^{-1}(U)$. Then $\overline{q}(d)$ is the desired element $d^{\prime}\in D^{\prime}$ with $d^{\prime}\in U$ and so $U$ is Scott-open. ∎
Remark 3.10.
Note that a similar proof shows this for any Scott-continuous order quotient between dcpos.
@@
The map $\overline{q}\colon\Sigma^{\Sigma^{G}}\to\mathcal{O}X$ is equal to $q\colon\Sigma^{\Sigma^{G}}\to\Sigma^{X}$ as defined in Definition 3.7 whenever $X$ is locally compact.
@@
We first note that the codomains agree since $\mathcal{O}X=\Sigma^{X}$ for locally compact $X$. Now recall that $q=\Sigma^{i_{X}}$. The natural transformation $(-)^{i_{X}}$ is the mate of ()×$(-)\times i_{X}$ with respect to the adjunction ()×$(-)\times X\dashv(-)^{X}$ and so we have the following commutative diagram.
{tikzpicture}
@@
The failure of $\mathcal{O}X$ with the Scott topology to be an exponential $\Sigma^{X}$ when $X$ is not locally compact is due to this quotient map being badly behaved in general. Indeed, if there were a section to $q$ as in Proposition 3.8 we could show $X$ to be locally compact after all. (The proofs of the following two results require us to assume $X$ itself is spatial, but are useful for intuition. We give pointfree results in Section 3.2.2.)
@@
Let $\mathrm{ev}=\widetilde{\mathrm{ev}}\circ(s\times X)$. On points we have $\mathrm{ev}(U,x)$ = $\widetilde{\mathrm{ev}}(s(U),x)$, which is equal to $\top$ precisely when $i_{X}(x)\in s(U)$ in $\Sigma^{G}$ and hence when $x\in U$ in $X$. Thus, $\mathrm{ev}$ is the evaluation map for exponential of the underlying sets. Moreover, $\mathrm{ev}$ is continuous, since it is a composition of continuous maps. Every continuous map from $Y\times X\to\Sigma$ factors through $\mathrm{ev}$ to give a set-theoretic function from $Y$ to $\Sigma^{X}$. But this map is continuous since $\mathcal{O}X$ has the final topology with respect to $\overline{q}$. Thus, $\mathcal{O}X$ is indeed the exponential $\Sigma^{X}$ in $\mathrm{Top}$. This is known to imply that $X$ is exponentiable in $\mathrm{Top}$, and even in $\mathrm{Loc}$. Hence, $X$ is locally compact. ∎
In fact, we can say more. It is known that lack of exponential objects is related to the fact that quotients are not stable under pullback (see [4]). We have the following following result.
@@
Here $\mathrm{ev}$ is defined as the evaluation map from the exponential of the underlying sets. It is continuous since the composite $\widetilde{\mathrm{ev}}=\mathrm{ev}\circ(q\times X)$ is continuous and $\overline{q}\times X$ is a quotient map.

### v4

@@ 3.1. Presentations and weak exponentials
Let $X$ and $B$ be objects in some category. We say $Z$ is a weak exponential with base $B$ and exponent $X$ if there exists a morphism $\widetilde{\mathrm{ev}}\colon Z\times X\to B$ such that for any $f\colon Y\times X\to B$ there exists a not necessarily unique $\overline{f}\colon Y\to Z$ such that the following diagram commutes.
@@
$A\times X$$A\times\Sigma^{G}$$\Sigma$$\Sigma^{\Sigma^{G}}\times\Sigma^{G}$$\Sigma^{\Sigma^{G}}\times X$$A\times i_{X}$$u^{\prime}$$\overline{u}^{\prime}\times X$$\overline{u}^{\prime}\times\Sigma^{G}$$\Sigma^{\Sigma^{G}}\times i_{X}$$\mathrm{ev}$$u$
Observe that $u$ defines an open in $A\times X$. Since $A\times X$ is a subspace of $A\times\Sigma^{G}$, there exists a (not necessarily unique) open $u^{\prime}$ in $A\times\Sigma^{G}$ which restricts to give $u$. By the universal property of the exponential $\Sigma^{\Sigma^{G}}$, we obtain a morphism $\overline{u}^{\prime}$ making the bottom right triangle commute. The left-hand trapezium commutes since both composites give $\overline{u}^{\prime}\times i_{X}$ and then a simple diagram chase confirms that $u=\widetilde{\mathrm{ev}}\circ(\overline{u}^{\prime}\times X)$. ∎
The above construction is summarised in the following definition.
@@
Let $X$ be a locale with $\mathcal{O}X=\langle G\mid R\rangle$. We define the associated (idealised) machine space to be $\Sigma^{\Sigma^{G}}$ together with the evaluation map $\widetilde{\mathrm{ev}}\colon\Sigma^{\Sigma^{G}}\times X\to\Sigma$ defined in Proposition˜3.2. ∎
We explore the intuition behind the concept of machine space in the following section.
3.2. Philosophical Interlude
In this section we examine the relationship between opens, i.e. verifiable properties, and the processes that semi-decide them. We view these processes as being carried out by machines, which we will identify with elements of the weak exponential $\Sigma^{\Sigma^{G}}$.
Given a space $X$ with $\mathcal{O}X=\langle G\mid R\rangle$ we interpret each generator $g\in G$ as a basic black-box machine which takes as input the points of $X$ and either halts after some finite amount of time, or runs forever. More complicated machines can be constructed from these basic machines. For instance, we may construct a composite machine by taking a finite collection of machines and insist that each machine halts on the given point before the composite machine is considered to have halted. We can also construct a further machine by running any number of these composite machines in parallel and halting on a point if at least one branch halts. This corresponds to taking formal joins of finite formal meets of generators and so we see that these machines are precisely the elements of the frame frame on $G$ – that is, the points of the machine space $\Sigma^{\Sigma^{G}}$.
Definition 3.4.
Notice that $\Sigma^{\Sigma^{G}}$ by itself contains no information about the behaviour of the machines on inputs from the locale $X$ with presentation $\langle G\mid R\rangle$, as this data is contained in the relation $R$. We can recover this information from the evaluation map $\widetilde{\mathrm{ev}}\colon\Sigma^{\Sigma^{G}}\times X\to\Sigma$ of Proposition˜3.2, which records whether a machine halts on a given point.
@@ 3.2. Philosophical Interlude
In fact, for many purposes any weak exponential with base $\Sigma$ exponent $X$ behaves like a machine space for $X$. This could be interesting to explore further. However, for this paper we restrict our consideration to $\Sigma^{\Sigma^{G}}$, since the compactness algorithm given in Section˜4 is dependent on the presentation of the locale by generators and relations.
A point $x\in X$ is accepted by a machine $m$ if there exists some $J_{k}$ such that $x$ is accepted by each ‘basic machine’ $g_{j}$ for all $j\in J_{k}$ (i.e. such that $x$ lies in the open given by $g_{j}$ for all such $j$). In this way, each machine $m$ will define an open on which it halts. Many of these machines will correspond to the same opens, as described by the relation $R$, so that each open in $X$ corresponds to an equivalence class of machines. In Definition˜3.9 of Section˜3.3 we will see that when the exponential $\Sigma^{X}$ exists there is a quotient map $\Sigma^{\Sigma^{G}}\twoheadrightarrow\Sigma^{X}$ assigning machines to opens.
Remark 3.6.
Recall that the failure of the exponential $\Sigma^{X}$ to always exist poses a puzzle for the verifiable interpretation of topology. Luckily, we have seen that machine spaces always exists. It is now helpful to consider in detail how such spaces might correspond to concrete objects that can be learnt about via physical measurements. Note that with such “real-world” models the set of generators is assumed to be countable.
Model 3.7.
Let $X$ be a space with $\mathcal{O}X=\langle G\mid R\rangle$. We take each generator $g\in G$ to be a black box which accepts points of $X$ as input and either halts after some finite time or runs forever. We imagine these generators arranged in a line. Additionally there are a number of mobile machines which themselves accept points of $X$ as input and will move from generator to generator testing the point in the following manner.
The machine progresses through different stages. On the $i^{\text{th}}$ stage it selects a group of finitely many generators, writes the number $i$ onto them and (if they are not already running) runs them with the point as input. The generators are left to run as the next stage is started. If at any point every generator in a group has halted, the machine halts.
Such a machine instantiates the element $\bigvee_{i\in I}\bigwedge_{j\in J_{i}}g_{j}$ where each $J_{i}$ is the finite set of generators selected at stage $i$.
To accommodate the fact that $\vee$ and $\wedge$ are commutative, absorptive and idempotent we require the following of our machines.
  1. (1)
    The order in which the machine visits each stage and the generators therein is chosen non-deterministically,
  2. (2)
    If it visits the generators $S\subseteq G$ in a stage, then in some other stage it must visit $T\supseteq S$, for all such finite $T\subseteq G$,
  3. (3)
    Finally, no two stages will have the exact same generators, and no stages will contain the same generator twice.
Note that associativity and distributivity are automatically satisfied with distributivity following from the fact that the formal operations are already reduced to the canonical form of a join of meets. These axioms together are then enough to give a frame ∎
What properties can we verify about the machines in this model? Of course, for a real implementation there are a number of irrelevant properties, such as the precise position of the machines, but taking an idealised perspective and viewing non-deterministic choices as conferring no additional information, there are still some nontrivial verifiable properties. This indicates that these machines are not complete black boxes and that there is some information that leaks while they run: we can semi-decide if there exists some parallel branch of a machine $m$ such that each of the generators that run in that branch are contained in $U$. This is exactly the open ${\boxtimes}U$. Recall that the opens of this form generate the frame $\mathcal{O}\Sigma^{\Sigma^{G}}$.
We can now answer the question posed in the introduction. While we have seen that machine space is reasonably concrete, the ‘space of opens’ is less so. In the flawed argument for the existence of a space of opens from Section˜1, one assumes that given an open and a point, it is possible to semi-decide membership of the point in the open. But an open merely describes the verifiable property under consideration — it is a machine that actually carries out this verification.
In the following subsection, we will show that the quotient map $\Sigma^{\Sigma^{G}}\twoheadrightarrow\mathcal{O}X$ has a section whenever the space $X$ is locally compact. From the above perspective, this fact perfectly explains why in case a continuous evaluation map exists for $\mathcal{O}X$: given an open $u$ and a point $x$, one can apply the section to $u$ to acquire a machine $s(u)$ which can then verify membership of $x$ in $u$.
On the other hand, when $X$ is not locally compact the argument breaks down and we will see that there is no way to continuously map opens to machines in this case.
Remark 3.8.
Note that since we not care which specific machine representing a given open is selected, we can relax the requirement of having an actual section to having a sufficiently well-behaved quotient. In particular, suplattice and preframe homomorphisms between frames are known to correspond to angelically and demonically nondeterministic maps respectively [24, 31, 30] and it would seem that a nondeterministic choice of representative for each open would be sufficient to construct the evaluation map. Indeed, open and proper quotients, or more generally, triquotients, have such ‘non-deterministic sections’ and these are strong enough conditions on the quotient map to entail that $X$ is exponentiable. We will discuss the role of nondeterminism in more detail in a later paper.
@@ 3.3. Relationship between machine space and the space of opens
$\Sigma^{\Sigma^{G}}\times X$$\Sigma^{X}\times X$$\Sigma$$\Sigma^{X}$$\Sigma^{\Sigma^{G}}$$q\times X$$\mathrm{ev}$$\widetilde{\mathrm{ev}}$$q$
@@
Let $X$ be a locally compact locale with $\mathcal{O}X=\langle G\mid R\rangle$. The map $q\colon\Sigma^{\Sigma^{G}}\to\Sigma^{X}$ as defined in Definition˜3.9 has a section $s$ satisfying $\widetilde{\mathrm{ev}}\circ(s\times X)=\mathrm{ev}$.
@@
$\Sigma^{\Sigma^{G}}\times X$$\Sigma^{X}\times X$$\Sigma$$\Sigma^{X}$$\Sigma^{\Sigma^{G}}$$q\times X$$s\times X$$\mathrm{ev}$$\widetilde{\mathrm{ev}}$qqss
It is immediate that $\widetilde{\mathrm{ev}}\circ(s\times X)=\mathrm{ev}$ and so all that remains is to prove that $s$ is a section of $q$.
By the diagram we have $\mathrm{ev}\circ(q\times X)\circ(s\times X)=\widetilde{\mathrm{ev}}\circ(s\times X)=\mathrm{ev}=\mathrm{ev}\circ(\mathrm{id}_{\Sigma^{X}}\times X)$. Hence by the uniqueness condition of the universal property of the ‘true’ exponential $\Sigma^{X}$, we have $q\circ s=\mathrm{id}_{\Sigma^{X}}$, as required. ∎
@@
Conversely, if opens can be continuously assigned to machines, then we expect the exponential to exist. In fact, as we will see, it suffices for the map $q$ to be a ‘sufficiently good’ quotient map. (For example, we might only have a ‘non-deterministic section’.)
Recall that if a locale $X$ is given by $\mathcal{O}X=\langle G\mid R\rangle$ then the frame $\mathcal{O}X$ is a quotient of $\mathcal{O}(\Sigma^{G})$ by the congruence generated from the relations $R$. Let us call this frame quotient map $\overline{q}$ for notational convenience, but as a function it is the same as $i_{X}^{*}$. Recall that $\mathcal{O}(\Sigma^{G})$ equipped with the Scott topology is precisely the machine space $\Sigma^{\Sigma^{G}}$. We can now use $\overline{q}$ to induce a topology on $\mathcal{O}X$. We will make use of the following result.
Lemma 3.11.
Let $L$ and $M$ be posets admitting all joins and let $q\colon L\twoheadrightarrow M$ be join-preserving surjection. Then if $L$ is equipped with the Scott-topology, the quotient topology induced on $M$ is also the Scott-topology.
Proof.
We need to show that $U\subseteq M$ is Scott-open if and only if $q^{-1}(U)$ is Scott-open.
Now suppose $q^{-1}(U)$ is Scott-open. Consider $a\in U$ and suppose $a\leq b$. Take $a^{\prime}\in q^{-1}(\{a\})$ and $b^{\prime}\in q^{-1}(\{b\})$. We may assume $b^{\prime}\geq a^{\prime}$ by joining $a^{\prime}$ if necessary. Then $a^{\prime}\in q^{-1}(U)$ and hence $b^{\prime}\in q^{-1}(U)$, since $q^{-1}(U)$ is upward closed. Thus, $b=q(b^{\prime})\in U$ and so $U$ is an upset. Next suppose $\operatorname*{\bigvee{}^{\!\uparrow}_{\hskip-0.60275pt\hskip 2.2604pt}}D\in U$. Set $D^{\prime}=q^{-1}({\downarrow}D)$. We note $D^{\prime}$ is directed. Certainly $D^{\prime}$ is nonempty. If $c,d\in D^{\prime}$ then there is an $e\in D$ with $q(c),q(d)\leq e$ by directedness of $D$. Thus, $q(c\vee d)=q(c)\vee q(d)\leq e\in{\downarrow}D$ and so $c\vee d\in D^{\prime}$. Furthermore, note that $q(D^{\prime})={\downarrow}D$ and $q(\operatorname*{\bigvee{}^{\!\uparrow}_{\hskip-0.60275pt\hskip 2.2604pt}}D^{\prime})=\operatorname*{\bigvee{}^{\!\uparrow}_{\hskip-0.60275pt\hskip 2.2604pt}}q(D^{\prime})=\operatorname*{\bigvee{}^{\!\uparrow}_{\hskip-0.60275pt\hskip 2.2604pt}}{\downarrow}D=\operatorname*{\bigvee{}^{\!\uparrow}_{\hskip-0.60275pt\hskip 2.2604pt}}D\in U$. Thus, $\operatorname*{\bigvee{}^{\!\uparrow}_{\hskip-0.60275pt\hskip 2.2604pt}}D^{\prime}\in q^{-1}(U)$ and so there is a $d^{\prime}\in D^{\prime}$ with $d^{\prime}\in q^{-1}(U)$, since $q^{-1}(U)$ is Scott-open. Now $q(d^{\prime})\in{\downarrow}D$ so that $q(d^{\prime})\leq d$ for some $d\in D$. Recalling that $U$ is an upset, we have $d\geq q(d^{\prime})\in U$ and so $d$ is the desired $d\in U$. Thus, $U$ is Scott-open. ∎
Corollary 3.12.
@@
The map $\overline{q}\colon\Sigma^{\Sigma^{G}}\to\mathcal{O}X$ is equal to $q\colon\Sigma^{\Sigma^{G}}\to\Sigma^{X}$ as defined in Definition˜3.9 whenever $X$ is locally compact.
@@
We first note that the codomains agree since $\mathcal{O}X=\Sigma^{X}$ for locally compact $X$. Now recall that $q=\Sigma^{i_{X}}$. The natural transformation $(-)^{i_{X}}$ is the mate of $(-)\times i_{X}$ with respect to the adjunction $(-)\times X\dashv(-)^{X}$ and so we have the following commutative diagram.
@@
The failure of $\mathcal{O}X$ with the Scott topology to be an exponential $\Sigma^{X}$ when $X$ is not locally compact is due to this quotient map being badly behaved in general. Indeed, if there were a section to $q$ as in Proposition˜3.10 we could show $X$ to be locally compact after all.
@@
We thank the reviewer for suggesting the following proof, which is simpler than the one we originally had.
@@
Since $\Sigma^{G}$ is locally compact, $\mathcal{O}(\Sigma^{G})$ is a continuous frame and in particular a continuous dcpo. But continuous dcpos are known to be stable under Scott-continuous retracts and so $\mathcal{O}X$ is also a continuous dcpo (see [2, Theorem 3.1.4.]). Thus, $\mathcal{O}X$ is a continuous frame and $X$ is locally compact.
To see that $\widetilde{\mathrm{ev}}\circ(s\times X)$ is the evaluation map, first note that, by Lemma˜3.13, $\overline{q}=q$ as defined in Definition˜3.9. So by the commutative diagram in Definition˜3.9, the evaluation map $\mathrm{ev}\colon\mathcal{O}X\times X\to\Sigma$ satisfies $\mathrm{ev}\circ(q\times X)=\widetilde{\mathrm{ev}}$. Hence, $\widetilde{\mathrm{ev}}\circ(s\times X)=\mathrm{ev}\circ(q\times X)\circ(s\times X)=\mathrm{ev}$. ∎
In fact, we can say more (see Remark˜3.8). It is known that the lack of exponential objects is related to the fact that quotients are not stable under products (see [6]). We have the following result. (This result requires us to assume $X$ is spatial and that we are in the category of topological spaces, though it is still useful for intuition. But see also Remark˜3.16.)
@@
Here $\mathrm{ev}$ is defined as the evaluation map from the exponential of the underlying sets. It is continuous since the composite $\widetilde{\mathrm{ev}}=\mathrm{ev}\circ(\overline{q}\times X)$ is continuous and $\overline{q}\times X$ is a quotient map.